In the table at right, how many workers are employed when marginal product reaches its maximum? Input of Labor (number Total Product (number of workers in weeks) of tablets produced OA 1 Ов.з Ос. 4 30 68 110 140 135
Answer #1
Input of labour Total product Marginal product
0 0 --
1 30 30-0=30
2 68 68-30=38
3 110 110-68=42
4 140 140-110=30
5 135 135-140=-5

Answer= b. 3 workers

At 3 workers, marginal product is maximum at 42.

If it helps, kindly upvote.